Session: Feb 2000 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K.Borkowski at Thu Feb 17 13:55:33 MET 2000) Success
The station file, as well as the .sum file, contained different time schedule then the .skd file since 6:49 or 6:55 UT. Of cource we have followed the .skd schedule. No other known problems.

Effelsberg (A. Kraus at Thu Feb 17 13:57:46 MET 2000) Minor failures
The observations was interrupted from UT0750-0850 because of snowfall.

Noto (C.Trigilio at Thu Feb 17 14:44:19 MET 2000) Minor failures
MK4 formatter used for this experiment. No signal in bbc because of wrong filter at the receiver -> lost first hour of observation. From 01:14 UT OK.

Westerbork (Tony Foley at Thu Feb 17 17:53:26 MET 2000) Failure
A confusion with which of two 21cm mixing schemes was to be used gave a loss of data from 00:05 to 09:00 UT. Only data from 09:00 to 12:00 will be any use. One tape (NOAA00016) was extremely dirty. Full tied array used.

Onsala (A. Polatidis at Mon Feb 21 12:49:15 MET 2000) Success

Jodrell (Alastair Gunn at Wed Feb 23 14:36:22 MET 2000) Success
No known problems.

Calibration (Alastair Gunn at Wed Apr 19 13:02:32 METDST 2000) Success
Calibration data available from vlbeer: see file feb00/eh007a.antab for details